The measure of an angle is 12 less than twice the measure of its supplement. What is the measure of the angle?

Answer by checkley79(3341)

X=2(180-X)-12
X=360-2X-12
X+2X=360-12
3X=348
X=348/3
X=116 DEGREES IS THE LARGER ANGLE.
180-116=64 DEGREES FOR THE SMALLER ANGLE.
PROOF:
116=2*64-12
116=128-12
116=116